OpenStudy (anonymous):

What are the coordinates of the center of the circle x^2+y^2-12x-6y-4=0

OpenStudy (anonymous):

i think it may -6,3 but not sure

OpenStudy (campbell_st):

do you know how to complete the square \[x^2 - 12x +y^2 - 6y = 4\]

OpenStudy (anonymous):

ik that the centre is: (-g, -f)

OpenStudy (campbell_st):

all you need to do to make certain is complete the square to both x and y \[x^2 -12x + (12/2)^2 + y^2 -6y + (6/2)^2 = 4 +(12/2)^2 + (6/2)^2\] factorising the terms in x and y, which are now perfect squares, will give the centre and radius